A bowl of water has a temperature of 50◦C. It is put into a refrigerator where the temperature is 5◦C. After 0.5 hours, the water is stirred and its temperature is measured to be 20◦C. It is then left to cool further, take t in Newton’s law of cooling to be measured in minutes. Use Newton’s law of cooling to predict when the temperature will be 10◦C.
1
Expert's answer
2020-06-28T18:01:52-0400

The Newton's law of cooling can be written in form

"T(t) = T_{\text{out}} + e^{-kt}(T_0-T_{\text{out}})." Here temperature is in Kelvins.

"293 = 278 + e^{-k\cdot 30}(323-278), \;\; e^{-30k} = \dfrac{15}{45}, \;\; k = 0.0366."

"283 = 278 + e^{-kt}(323-278), \;\; e^{-kt} = \dfrac{5}{45}, \;\; t =60\,\mathrm{m} = 1\,\mathrm{h}."